OpenStudy (anonymous):

Can somone factor x^2-6x-27?

OpenStudy (anonymous):

yes but you should learn how to do it?

OpenStudy (anonymous):

yes someone can

OpenStudy (anonymous):

It helps me understand if I see someone do it.

jhonyy9 (jhonyy9):

x*x -9x+3x -9*3 = x(x-9) +3(x-9) = (x+3)(x-9) ok ?

OpenStudy (anonymous):

Yeah! Thanks a lot.

jhonyy9 (jhonyy9):

good luck bye